# The Power of Iterative Filtering for Supervised Learning with (Heavy) Contamination

### v1

@@ 4.3 Tolerant Testable Learning
Previous work by [54] showed that the existence of $\mathcal{L}_{2}$ sandwiching polynomials with bounded coefficients, i.e., polynomials $p_{\mathrm{up}},p_{\mathrm{down}}$ such that $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}^{*}}[(p_{\mathrm{up}}(\mathbf{x})-p_{\mathrm{down}}(\mathbf{x}))^{2}]\leq\epsilon$ implies tolerant testable learning. Here, we relax this condition to $\mathcal{L}_{1}$ sandwiching polynomials (Definition 4.3), without requiring any bound on the coefficients. Note that even for non-tolerant testable learning, where $\mathcal{L}_{1}$ sandwiching is known to suffice [48], all prior work required bounds on the coefficients.
@@
The following corollary follows from Theorem 4.8, combined with the fact that log-concave distributions are hypercontractive (see [10]), as well as the results of [73, 67] on the sandwiching degree of functions of halfspaces (see Section C.2). This is the first result for testably learning even a single halfspace with respect to log-concave distributions up to optimal error.
@@
We will now prove Theorem 4.8, using once more the iterative polynomial filtering algorithm of Theorem 3.2.
Proof of Theorem 4.8.
@@
    First, the algorithm runs the filtering procedure of Theorem 3.2 (that is, Algorithm 1) on input $(S_{\mathrm{inp}},S_{\mathrm{ref}},m=|S_{\mathrm{inp}}|,\ell,R=\frac{4\tau}{\epsilon}+{2},\epsilon/8)$ to form the filtered dataset $\bar{S}_{\mathrm{filt}}$.
@@
    The algorithm outputs $h(\mathbf{x})=\operatorname{sign}(\widehat{p}(\mathbf{x})+\widehat{\tau})$, where $\widehat{\tau}\in{\mathbb{R}}$ minimizes the one-dimensional objective $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{filt}}}[y\neq\operatorname{sign}(\widehat{p}(\mathbf{x})+\tau)]$ over $\tau\in{\mathbb{R}}$.
@@
It suffices to show that $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{filt}}}[y\neq h(\mathbf{x})]\leq\frac{m}{|\bar{S}_{\mathrm{filt}}|}(\min_{f\in\mathcal{C}}\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{\mathcal{D}}}[y\neq h(\mathbf{x})]+\epsilon/4)$. Let $f^{*}$ be the function that minimizes the error on $\bar{\mathcal{D}}$ and let $p_{\mathrm{up}},p_{\mathrm{down}}$ be its $\frac{\epsilon^{2}}{128\tau+64\epsilon}$-sandwiching polynomials. Then according to part 2 of Theorem 3.2, and since for $p=p_{\mathrm{up}}-p_{\mathrm{down}}$ we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[p(\mathbf{x})]\leq\frac{\epsilon^{2}}{128\tau+64\epsilon}$, we obtain:
@@
Therefore, we also have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{filt}}}[|f^{*}(\mathbf{x})-p_{\mathrm{down}}(\mathbf{x})|]\leq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{filt}}}[p_{\mathrm{up}}(\mathbf{x})-p_{\mathrm{down}}(\mathbf{x})]\leq\frac{\epsilon m}{8|S_{\mathrm{filt}}|}$. Observe, now that $h$ is the output of the low-degree polynomial regression algorithm and, hence, according to Theorem A.3 by [68], we have the following.
@@
Note that the distribution of each $\mathbf{z}^{(i)}$ is $\mathcal{D}^{*}$ and they are drawn independently. Due to a Chernoff bound, since $m$ is large enough, we have that, with high probability, $|S_{\mathrm{inp}}\setminus S_{\mathrm{cln}}|\leq m(\tau+\epsilon/8)$. According to Theorem 3.2, we have the following, with high probability:
@@ 5.1 Bounded Contamination
Furthermore, many of the lower bounds for agnostic learning seem to exclusively rule out algorithms with error guarantees of $\mathsf{opt}+\epsilon$ [37]. Therefore, these lower bounds do not transfer directly to learning with bounded contamination as defined in Definition 1.1. One natural way to go around this, is to observe that our result in Theorem 4.2 would also hold for the following stronger version of Definition 1.1.
@@
Note that Definition 5.1 is a generalization of agnostic learning, since one can obtain agnostic learning by setting $\eta=0$. Our Theorem 4.2 holds even under this definition, and all the lower bounds for agnostic learning (Table 4) are also inherited. For example, our results indicate that a runtime of $d^{\operatorname{poly}(\frac{1}{\epsilon},\log k)}$ is both sufficient and necessary for learning intersections of $k$ halfspaces in the agnostic BC setting. However, agnostic learning is known to be characterized by the $\mathcal{L}_{1}$-approximate degree [68, 42, 37], whereas we require the stronger assumption of low $\mathcal{L}_{2}$-approximate degree. Therefore, while for most interesting classes our results are essentially tight, a complete characterization of efficient learnability in the setting of Definition 5.1 remains open.
@@ 5.2 Heavy Contamination
We provide an information-theoretic lower bound for learning monotone functions with heavy contamination. Our lower bound highlights a separation between bounded and heavy contamination and justifies the need for a stronger notion of approximation than standard polynomial approximators. In particular, monotone functions over {±$\{\pm 1\}^{d}$ are known to admit $\epsilon$-approximate polynomials of degree $\ell(\epsilon)=\sqrt{d}/\epsilon$, and hence admit a BC-learner that runs in time $2^{{O}(\sqrt{d})}$ for $\epsilon=1/4$, but we show that every HC learner requires $2^{\Omega(d)}$ samples (and runtime).
@@
Let $\mathcal{C}$ be the class of monotone functions over {±$\{\pm 1\}^{d}$. Any HC-learner for $\mathcal{C}$ with respect to the uniform distribution over the hypercube $\operatorname{Unif}(\{\pm 1\}^{d})$ requires sample complexity $2^{\Omega(d)}$, even when $\epsilon=1/4$, $Q=2$ and $\delta=1/3$.
Proof of Theorem 5.2.
Suppose that the ground truth is a constant function, i.e., either $f^{*}\equiv 1$ or $f^{*}\equiv-1$, and that $m\leq 2^{d/10}/10$. Then, the adversary receives the $m$ clean examples $\bar{S}_{\mathrm{inp}}$ and draws $m$ additional i.i.d. examples $S^{\prime}$ from $\operatorname{Unif}(\{\pm 1\}^{d})$, labeling them according to $-f^{*}$. The input dataset is $\bar{S}_{\mathrm{inp}}=\bar{S}_{\mathrm{cln}}\cup\bar{S}^{\prime}$.
@@
We have a set $\bar{S}_{\mathrm{inp}}$ of $2m$ examples. It suffices to show that no pair of examples in $S_{\mathrm{inp}}$ is comparable, because, then, any labeling of these examples is consistent with some monotone function and, hence, $\mathsf{opt}_{\mathrm{total}}=0$. We overall have at most $4m^{2}$ pairs and each of them is comparable with probability at most $2\cdot(3/4)^{d}$. Therefore, by a union bound, the probability that there is a pair that is comparable is at most $108m^{2}(3/4)^{d}\leq 1/$ 10. ∎
@@ 6 Discussion
    Improved Error Guarantees: Our algorithms obtain error guarantees that are proven to be optimal in general (see Proposition B.3). However, it is an open question whether one can achieve error $\eta+\epsilon$ for BC-learning or error $\frac{1}{2}(Q\cdot\mathsf{opt}_{\mathrm{total}}+\mathsf{opt}_{\mathrm{clean}})+\epsilon$ for HC-learning for special classes and distributions, perhaps by allowing the learner to output randomized hypotheses (also known as probabilistic concepts).
@@
    Heavy Contamination without Label Noise: Our definition for learning with heavy contamination allows the adversary to add arbitrary labeled examples, as long as the resulting contaminated dataset can be approximately labeled by some classifier in the learned class up to a small number of mistakes. Note that the optimal classifier can also be chosen adversarially, in the sense that we do not require that the optimal classifier for the contaminated dataset be the same as the optimal classifier on the clean distribution. This property is crucially used in our lower bound for learning monotone functions with heavy contamination (Theorem 5.2): the hard instance is perfectly labeled by some monotone function, but this monotone function is far from the one that generates the clean labels.
    On the other hand, even in the realizable HC setting—where the clean labels are given by some concept $f^{*}\in\mathcal{C}$ and the adversary is required to add points of the form $(\mathbf{x},f^{*}(\mathbf{x}))$—it is not clear when efficient learning is possible. This is because for many classes, the assumptions on the distribution of $\mathbf{x}$ are crucial, even in the absence of label noise. It is therefore an interesting open question whether improved HC-learning guarantees can be achieved in the realizable setting.
@@ Appendix A Additional Tools
$\frac{1}{2}\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[(p(\mathbf{x}))^{2}]\leq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(p(\mathbf{x}))^{2}]\leq 2\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[(p(\mathbf{x}))^{2}]\,,\,\text{ for any }p\text{ of degree at most }\ell$
@@
Let $\mathcal{D}$ be some ${A}$-hypercontractive distribution in $d$ dimensions for some $A\geq 1$. Let $S$ be a set of $m$ i.i.d. samples from $\mathcal{D}$ and $\mathcal{P}_{\beta}(S)$ be the family of polynomials of degree at most $\ell$ such that $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(p(\mathbf{x}))^{2}]\leq\beta$, where $\ell,\beta,d\geq 1$. Then, for any $\epsilon,\delta\in(0,1)$, if $m\geq(CAd)^{2\ell}(\log\frac{1}{\delta})^{4\ell+1}$, and $B\geq\sqrt{{2\beta}/{\epsilon}}\cdot(d+1)^{\ell/2}$ the following holds with probability at least $1-\delta$ over $S$:
@@
Proof of Lemma A.2.
We first apply Lemma A.1 to show that, with probability at least $1-\delta$ over $S$, for any $p\in\mathcal{P}_{\beta}(S)$ we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[(p(\mathbf{x}))^{2}]\leq 2\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(p(\mathbf{x}))^{2}]\leq 2\beta$.
@@
Let $\bar{S}$ be a set of labeled examples and $\mathcal{C}$ be a concept class such that for each $f\in\mathcal{C}$ there is some polynomial $p$ of degree at most $\ell$ such that $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[|f(\mathbf{x})-p(\mathbf{x})|]\leq\epsilon$. Then, the $-\ell$ $\mathcal{L}_{1}$ polynomial regression algorithm of [68] outputs, in time $\operatorname{poly}(\lvert\bar{S}\rvert d^{\ell}/\epsilon)$, a $-\ell$ polynomial threshold function $h$ such that $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq h(\mathbf{x})]\leq\min_{f\in\mathcal{C}}\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq f(\mathbf{x})]+\epsilon$.
We also give the definition of $\mathcal{L}_{1}$ approximating polynomials, which are sufficient for agnostic learning, but it is not clear whether they suffice for BC-learning. We instead use the notion of $\mathcal{L}_{2}$ approximating polynomials (Definition 4.1).
@@ Appendix B Error Lower Bounds for HC-Learning
An algorithm $\cal M$ is called a (randomized) $Q$-HC strategy if, upon receiving a labeled set $\bar{S}_{\mathrm{cln}}$, it outputs a set $\bar{S}_{\mathrm{inp}}$ of size $|\bar{S}_{\mathrm{inp}}|\leq Q|\bar{S}_{\mathrm{cln}}|$ such that $\bar{S}_{\mathrm{inp}}\subseteq\bar{S}_{\mathrm{cln}}$. For $m\in\mathbb{N}$ and a labeled distribution $\bar{\mathcal{D}}$, we denote with ${\cal M}(\bar{\mathcal{D}}^{m})$ the distribution of the output $\bar{S}_{\mathrm{inp}}$ of $\cal M$ on a set of $m$ i.i.d. samples $\bar{S}_{\mathrm{cln}}$ from $\bar{\mathcal{D}}$. Note that $\bar{S}_{\mathrm{inp}}$ depends on the random choice of $\bar{S}_{\mathrm{cln}}$, as well as the randomness of $\cal M$.
@@
Let $\mathcal{C}$ be any concept class such that there is some $f\in\mathcal{C}$ for which $-f\in\mathcal{C}$. For any $Q\in\{2,3,\dots\}$, any unlabeled distribution $\mathcal{D}^{*}$, any $\rho\in[0,1/2)$, $\epsilon,\delta>0$, and any $m\geq\frac{\log(2/\delta)}{2\epsilon^{2}}$, there is a $Q$-HC strategy $\cal M$ and a distribution $\bar{\mathcal{D}}$ whose marginal is $\mathcal{D}^{*}$ such that, if we let $\bar{\mathcal{D}}^{\prime}$ be the distribution of $(\mathbf{x},-y)$, where $(\mathbf{x},y)\sim\bar{\mathcal{D}}$, then the following hold.
@@
    If $m_{1}\geq m_{2}$, then the adversary adds $M-m$ examples drawn by $\mathcal{D}$ and labeled by $-f(\mathbf{x})$.
@@
    If $m_{1}<m_{2}$, then the adversary adds $m_{2}-m_{1}$ examples drawn by $\mathcal{D}$ and labeled by $f(\mathbf{x})$, as well as $M-2m_{2}$ examples drawn by $\mathcal{D}$ and labeled by $-f(\mathbf{x})$.
@@
Given any realization $m_{i}$ of $\xi$ with $m_{i}>m/2$, we can equivalently form the input examples $\bar{S}_{\mathrm{inp}}$ by drawing $M$ i.i.d. samples from $\mathcal{D}$ and labeling $\xi$ of them according to $f(\mathbf{x})$ and $M-\xi$ of them according to $-f(\mathbf{x})$. We have that ${\cal M}(\bar{\mathcal{D}}_{1}^{m})|_{\xi>m/2}={\cal M}(\bar{\mathcal{D}}_{2}^{m})|_{\xi>m/2}$.
Given any realization $m_{i}$ of $\xi$ with $m_{i}\leq m/2$, we can equivalently form the input examples $\bar{S}_{\mathrm{inp}}$ by drawing $M$ i.i.d. samples from $\mathcal{D}$ and labeling $m-\xi$ of them according to $f(\mathbf{x})$ and $M-m+\xi$ of them according to $-f(\mathbf{x})$. We, therefore, have ${\cal M}(\bar{\mathcal{D}}_{1}^{m})|_{\xi\leq m/2}={\cal M}(\bar{\mathcal{D}}_{2}^{m})|_{\xi\leq m/2}$.
@@
For finite feature spaces and the particular choice of the class of all functions, we obtain a sharper information-theoretic lower bound on the error. Note that this lower bound captures the problem of HC-learning halfspaces over {±$\{\pm 1\}$ in one dimension.
@@
Suppose that $|\mathcal{X}|<\infty$ and $\mathcal{C}$ is any concept class from $\mathcal{X}$ to {±$\{\pm 1\}$. Then, on any value of $Q$ in $\{2,3,\dots\}$, any (distribution-specific) HC-learner will output a hypothesis with error $1$ with probability at least $2^{-|\mathcal{X}|}$ on some instance where $Q\cdot\mathsf{opt}_{\mathrm{total}}=1$. Moreover, if $\mathcal{C}=\{\pm 1\}^{\mathcal{X}}$, then the instance satisfies $\mathsf{opt}_{\mathrm{clean}}=0$.
Proof of Proposition B.3.
@@
Note that the input dataset is completely independent from the clean labels, since $\bar{S}_{\mathrm{cln}}\cup\bar{S}_{\mathrm{cln}}^{\prime}$ can be constructed equivalently by choosing an arbitrary function $g:\mathcal{X}\to\{\pm 1\}$ to label the points in $\bar{S}_{\mathrm{cln}}$ and then label the points in $\bar{S}_{\mathrm{cln}}^{\prime}$ according to $-g$.
Therefore, we may assume that the ground truth labels are generated by the function $f^{*}$, where $-f^{*}$ is the most likely output of the HC-learner on a dataset of the form $\bar{S}_{\mathrm{cln}}\cup\bar{S}_{\mathrm{cln}}^{\prime}\cup\bar{S}^{\prime}$. Since there are at most $2^{|\mathcal{X}|}$ possible functions, $\mathcal{A}$ must output $-f^{*}$ with probability at least $2^{-|\mathcal{X}|}$. When $\mathcal{A}$ outputs $-f^{*}$, the error is $1$.

### v2

@@ 4.3 Tolerant Testable Learning
Previous work by [49] showed that the existence of $\mathcal{L}_{2}$ sandwiching polynomials with bounded coefficients, i.e., polynomials $p_{\mathrm{up}},p_{\mathrm{down}}$ such that $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}^{*}}[(p_{\mathrm{up}}(\mathbf{x})-p_{\mathrm{down}}(\mathbf{x}))^{2}]\leq\epsilon$ implies tolerant testable learning. Here, we relax this condition to $\mathcal{L}_{1}$ sandwiching polynomials (Definition˜4.3), without requiring any bound on the coefficients. Note that even for non-tolerant testable learning, where $\mathcal{L}_{1}$ sandwiching is known to suffice [51], all prior work required bounds on the coefficients.
@@
The following corollary follows from Theorem˜4.8, combined with the fact that log-concave distributions are hypercontractive (see [11]), as well as the results of [73, 62] on the sandwiching degree of functions of halfspaces (see Section˜C.2). This is the first result for testably learning even a single halfspace with respect to log-concave distributions up to optimal error.
@@
We will now prove Theorem˜4.8, using once more the iterative polynomial filtering algorithm of Theorem˜3.2.
Proof of Theorem˜4.8.
@@
    First, the algorithm runs the filtering procedure of Theorem˜3.2 (that is, Algorithm˜1) on input $(S_{\mathrm{inp}},S_{\mathrm{ref}},m=|S_{\mathrm{inp}}|,\ell,R=\frac{4\tau}{\epsilon}+{2},\epsilon/8)$ to form the filtered dataset $\bar{S}_{\mathrm{filt}}$.
@@
[…]
@@
It suffices to show that $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}_{\mathrm{filt}}}[y\neq h(\mathbf{x})]\leq\frac{m}{|\bar{S}_{\mathrm{filt}}|}(\min_{f\in\mathcal{C}}\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{\mathcal{D}}}[y\neq h(\mathbf{x})]+\epsilon/4)$. Let $f^{*}$ be the function that minimizes the error on $\bar{\mathcal{D}}$ and let $p_{\mathrm{up}},p_{\mathrm{down}}$ be its $\frac{\epsilon^{2}}{128\tau+64\epsilon}$-sandwiching polynomials. Then according to part 2 of Theorem˜3.2, and since for $p=p_{\mathrm{up}}-p_{\mathrm{down}}$ we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[p(\mathbf{x})]\leq\frac{\epsilon^{2}}{128\tau+64\epsilon}$, we obtain:
@@
Therefore, we also have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{filt}}}[|f^{*}(\mathbf{x})-p_{\mathrm{down}}(\mathbf{x})|]\leq\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S_{\mathrm{filt}}}[p_{\mathrm{up}}(\mathbf{x})-p_{\mathrm{down}}(\mathbf{x})]\leq\frac{\epsilon m}{8|S_{\mathrm{filt}}|}$. Observe, now that $h$ is the output of the low-degree polynomial regression algorithm and, hence, according to Theorem˜A.3 by [60], we have the following.
@@
Note that the distribution of each $\mathbf{z}^{(i)}$ is $\mathcal{D}^{*}$ and they are drawn independently. Due to a Chernoff bound, since $m$ is large enough, we have that, with high probability, $|S_{\mathrm{inp}}\setminus S_{\mathrm{cln}}|\leq m(\tau+\epsilon/8)$. According to Theorem˜3.2, we have the following, with high probability:
@@ 5.1 Bounded Contamination
Furthermore, many of the lower bounds for agnostic learning seem to exclusively rule out algorithms with error guarantees of $\mathsf{opt}+\epsilon$ [35]. Therefore, these lower bounds do not transfer directly to learning with bounded contamination as defined in Definition˜1.1. One natural way to go around this, is to observe that our result in Theorem˜4.2 would also hold for the following stronger version of Definition˜1.1.
@@
Note that Definition˜5.1 is a generalization of agnostic learning, since one can obtain agnostic learning by setting $\eta=0$. Our Theorem˜4.2 holds even under this definition, and all the lower bounds for agnostic learning (Table˜4) are also inherited. For example, our results indicate that a runtime of $d^{\operatorname{poly}(\frac{1}{\epsilon},\log k)}$ is both sufficient and necessary for learning intersections of $k$ halfspaces in the agnostic BC setting. However, agnostic learning is known to be characterized by the $\mathcal{L}_{1}$-approximate degree [60, 22, 35], whereas we require the stronger assumption of low $\mathcal{L}_{2}$-approximate degree. Therefore, while for most interesting classes our results are essentially tight, a complete characterization of efficient learnability in the setting of Definition˜5.1 remains open.
@@ 5.2 Heavy Contamination
We provide an information-theoretic lower bound for learning monotone functions with heavy contamination. Our lower bound highlights a separation between bounded and heavy contamination and justifies the need for a stronger notion of approximation than standard polynomial approximators. In particular, monotone functions over $\{\pm 1\}^{d}$ are known to admit $\epsilon$-approximate polynomials of degree $\ell(\epsilon)=\sqrt{d}/\epsilon$, and hence admit a BC-learner that runs in time $2^{{O}(\sqrt{d})}$ for $\epsilon=1/4$, but we show that every HC learner requires $2^{\Omega(d)}$ samples (and runtime).
@@
Let $\mathcal{C}$ be the class of monotone functions over $\{\pm 1\}^{d}$. Any HC-learner for $\mathcal{C}$ with respect to the uniform distribution over the hypercube $\operatorname{Unif}(\{\pm 1\}^{d})$ requires sample complexity $2^{\Omega(d)}$, even when $\epsilon=1/4$, $Q=2$ and $\delta=1/3$.
Proof of Theorem˜5.2.
Suppose that the ground truth is a constant function, i.e., either $f^{*}\equiv 1$ or $f^{*}\equiv-1$, and that $m\leq 2^{d/10}/10$. Then, the adversary receives the $m$ clean examples $\bar{S}_{\mathrm{inp}}$ and draws $m$ additional i.i.d. examples $S^{\prime}$ from $\operatorname{Unif}(\{\pm 1\}^{d})$, labeling them according to $f^{*}$. The input dataset is $\bar{S}_{\mathrm{inp}}=\bar{S}_{\mathrm{cln}}\cup\bar{S}^{\prime}$.
@@
We have a set $\bar{S}_{\mathrm{inp}}$ of $2m$ examples. It suffices to show that no pair of examples in $S_{\mathrm{inp}}$ is comparable, because, then, any labeling of these examples is consistent with some monotone function and, hence, $\mathsf{opt}_{\mathrm{total}}=0$. We overall have at most $4m^{2}$ pairs and each of them is comparable with probability at most $2\cdot(3/4)^{d}$. Therefore, by a union bound, the probability that there is a pair that is comparable is at most $8m^{2}(3/4)^{d}\leq 1/10$. ∎
@@ 6 Discussion
    Improved Error Guarantees: Our algorithms obtain error guarantees that are proven to be optimal in general (see Proposition˜B.3). However, it is an open question whether one can achieve error $\eta+\epsilon$ for BC-learning or error $\frac{1}{2}(Q\cdot\mathsf{opt}_{\mathrm{total}}+\mathsf{opt}_{\mathrm{clean}})+\epsilon$ for HC-learning for special classes and distributions, perhaps by allowing the learner to output randomized hypotheses (also known as probabilistic concepts).
@@
    Heavy Contamination without Label Noise: Our definition for learning with heavy contamination allows the adversary to add arbitrary labeled examples, as long as the resulting contaminated dataset can be approximately labeled by some classifier in the learned class up to a small number of mistakes. Note that the optimal classifier can also be chosen adversarially, in the sense that we do not require that the optimal classifier for the contaminated dataset be the same as the optimal classifier on the clean distribution. This property is crucially used in our lower bound for learning monotone functions with heavy contamination (Theorem˜5.2): the hard instance is perfectly labeled by some monotone function, but this monotone function is far from the one that generates the clean labels.
    On the other hand, even in the realizable HC setting—where the clean labels are given by some concept $f^{*}\in\mathcal{C}$ and the adversary is required to add points of the form $(\mathbf{x},f^{*}(\mathbf{x}))$—it is not clear when efficient learning is possible. This is because for many classes, the assumptions on the distribution of $\mathbf{x}$ are crucial, even in the absence of label noise. Note that this setting is closer in spirit to prior work on semi-random analysis [71], where the adversary is monotone: from an information-theoretic perspective, the added examples can only help.
    One concrete question is whether monotone functions can be learned in time $2^{\tilde{O}_{\epsilon}(\sqrt{d})}$ under realizable heavy contamination, since our lower bound does not hold in this setting.
@@ Appendix A Additional Tools
[…]
@@
Let $\mathcal{D}$ be some A{A}-hypercontractive distribution in $d$ dimensions for some $A\geq 1$. Let $S$ be a set of $m$ i.i.d. samples from $\mathcal{D}$ and $\mathcal{P}_{\beta}(S)$ be the family of polynomials of degree at most $\ell$ such that $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(p(\mathbf{x}))^{2}]\leq\beta$, where $\ell,\beta,d\geq 1$. Then, for any $\epsilon,\delta\in(0,1)$, if $m\geq(CAd)^{2\ell}(\log\frac{1}{\delta})^{4\ell+1}$, and $B\geq\sqrt{{2\beta}/{\epsilon}}\cdot(d+1)^{\ell/2}$ the following holds with probability at least $1-\delta$ over $S$:
@@
Proof of Lemma˜A.2.
We first apply Lemma˜A.1 to show that, with probability at least $1-\delta$ over $S$, for any $p\in\mathcal{P}_{\beta}(S)$ we have $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim\mathcal{D}}[(p(\mathbf{x}))^{2}]\leq 2\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[(p(\mathbf{x}))^{2}]\leq 2\beta$.
@@
Let $\bar{S}$ be a set of labeled examples and $\mathcal{C}$ be a concept class such that for each $f\in\mathcal{C}$ there is some polynomial $p$ of degree at most $\ell$ such that $\operatorname*{\mathbb{E}}_{\mathbf{x}\sim S}[|f(\mathbf{x})-p(\mathbf{x})|]\leq\epsilon$. Then, the degree-$\ell$ $\mathcal{L}_{1}$ polynomial regression algorithm of [60] outputs, in time $\operatorname{poly}(\lvert\bar{S}\rvert d^{\ell}/\epsilon)$, a degree-$\ell$ polynomial threshold function $h$ such that $\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq h(\mathbf{x})]\leq\min_{f\in\mathcal{C}}\operatorname*{\mathbb{P}}_{(\mathbf{x},y)\sim\bar{S}}[y\neq f(\mathbf{x})]+\epsilon$.
We also give the definition of $\mathcal{L}_{1}$ approximating polynomials, which are sufficient for agnostic learning, but it is not clear whether they suffice for BC-learning. We instead use the notion of $\mathcal{L}_{2}$ approximating polynomials (Definition˜4.1).
@@ Appendix B Error Lower Bounds for HC-Learning
[…]
@@
[…]
@@
    If $m_{1}\geq m_{2}$, then the adversary adds MmM-m examples drawn by $\mathcal{D}$ and labeled by $f(\mathbf{x})$.
@@
    If $m_{1}<m_{2}$, then the adversary adds $m_{2}-m_{1}$ examples drawn by $\mathcal{D}$ and labeled by $f(\mathbf{x})$, as well as $M-2m_{2}$ examples drawn by $\mathcal{D}$ and labeled by $f(\mathbf{x})$.
@@
Given any realization $m_{i}$ of $\xi$ with $m_{i}>m/2$, we can equivalently form the input examples $\bar{S}_{\mathrm{inp}}$ by drawing $M$ i.i.d. samples from $\mathcal{D}$ and labeling $\xi$ of them according to $f(\mathbf{x})$ and $M-\xi$ of them according to $f(\mathbf{x})$. We have that ${\cal M}(\bar{\mathcal{D}}_{1}^{m})|_{\xi>m/2}={\cal M}(\bar{\mathcal{D}}_{2}^{m})|_{\xi>m/2}$.
Given any realization $m_{i}$ of $\xi$ with $m_{i}\leq m/2$, we can equivalently form the input examples $\bar{S}_{\mathrm{inp}}$ by drawing $M$ i.i.d. samples from $\mathcal{D}$ and labeling $m-\xi$ of them according to $f(\mathbf{x})$ and $M-m+\xi$ of them according to $f(\mathbf{x})$. We, therefore, have ${\cal M}(\bar{\mathcal{D}}_{1}^{m})|_{\xi\leq m/2}={\cal M}(\bar{\mathcal{D}}_{2}^{m})|_{\xi\leq m/2}$.
@@
For finite feature spaces and the particular choice of the class of all functions, we obtain a sharper information-theoretic lower bound on the error. Note that this lower bound captures the problem of HC-learning halfspaces over $\{\pm 1\}$ in one dimension.
@@
Suppose that $|\mathcal{X}|<\infty$ and $\mathcal{C}$ is any concept class from $\mathcal{X}$ to $\{\pm 1\}$. Then, on any value of $Q$ in $\{2,3,\dots\}$, any (distribution-specific) HC-learner will output a hypothesis with error $1$ with probability at least $2^{-|\mathcal{X}|}$ on some instance where $Q\cdot\mathsf{opt}_{\mathrm{total}}=1$. Moreover, if $\mathcal{C}=\{\pm 1\}^{\mathcal{X}}$, then the instance satisfies $\mathsf{opt}_{\mathrm{clean}}=0$.
Proof of Proposition˜B.3.
@@
Note that the input dataset is completely independent from the clean labels, since $\bar{S}_{\mathrm{cln}}\cup\bar{S}_{\mathrm{cln}}^{\prime}$ can be constructed equivalently by choosing an arbitrary function $g:\mathcal{X}\to\{\pm 1\}$ to label the points in $\bar{S}_{\mathrm{cln}}$ and then label the points in $\bar{S}_{\mathrm{cln}}^{\prime}$ according to g-g.
Therefore, we may assume that the ground truth labels are generated by the function $f^{*}$, where $f^{*}$ is the most likely output of the HC-learner on a dataset of the form $\bar{S}_{\mathrm{cln}}\cup\bar{S}_{\mathrm{cln}}^{\prime}\cup\bar{S}^{\prime}$. Since there are at most $2^{|\mathcal{X}|}$ possible functions, $\mathcal{A}$ must output $f^{*}$ with probability at least $2^{-|\mathcal{X}|}$. When $\mathcal{A}$ outputs $f^{*}$, the error is $1$.